You have been given the results of a health assessment screening that has been conducted by your medical staff on patients being treated for Type 2 diabetes at your clinic. The results from this standardized test, which scores individuals on a scale of 0 to 1000, form a normal distribution with a mean score of 680 and a standard deviation of 50.

To compare the results from patients of interest to a similar baseline test conducted by the Mayo Clinic, you will be required to convert their individual scores to match a normal distribution with a mean of 450 and a standard deviation of 25.

One of your patients has scored 700 on your local test. Convert this score to one that can be used to compare to the baseline test from the Mayo Clinic. Round your answer to two decimals.
Answer #1

Ans:

For local test:

z=(700-680)/50

z=0.4

Now, z score would be same for both distributions.

So,

required score,x=450+0.4*25=460